# On the minimax optimality of Flow Matching through the connection to kernel density estimation

### v1

@@ 4.2 Proofs and extended results of Section 3
By assumption $\varphi$ decays faster than $e^{-x}.$ Using the upper incomplete $\Gamma$-function as defined in Lemma A.2 and Lemma A.2, we obtain for $a>2$
@@
Next we want to choose $a$ in Lemma 4.1 and a neural network of finite size (depending on $n$) such that the bound of $\mathsf{W}_{1}(\mathbb{P}^{\psi^{n}(Z)},\mathbb{P}^{\hat{\psi}(Z)})$ decays at a rate of $n^{-\frac{1+\alpha}{2\alpha+d}}.$ Hence we need to approximate
and simultaneously control the Lipschitz constant of the approximation $\hat{v}$ on $(-a,a)^{d}.$ This is indeed sufficient, since using the same reasoning as in the proof of Lemma 4.1, it holds that if $x\in(-a,a)^{d}$, then $v_{t}^{n}(x)\in(-a,a)^{d}$ for every $t$. As $\hat{v}$ approximates $v^{n}$ in the supremum norm, the Lipschitz constant of $\hat{v}$ will be at least of the order of the Lipschitz constant of $v^{n}$. The same applies to the supremum norm. Hence we need bounds for both, the Lipschitz constant and the supremum norm of $v^{n}.$
Lemma 4.2.
Proof.
@@
$=\frac{t}{\sigma_{t}^{2}}\Big{(}\frac{\sum_{i=1}^{n}X_{ik}X_{i\ell}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{i})}-\frac{\big{(}\sum_{i=1}^{n}X_{ik}p_{t}(x|X_{i})\big{)}\big{(}\sum_{i=1}^{n}X_{i\ell}p_{t}(x|X_{i})\big{)}}{\big{(}\sum_{j=1}^{n}p_{t}(x|X_{i})\big{)}^{2}}\Big{)}.$
@@
From [Gühring et al., 2020, Corollary 4.2] we know that for any $f\in C^{2}$ with bounded second partial derivatives we can approximate $f$ and its Lipschitz constant simultaneously using a ReLU network as defined in (16). To be more precise, for every $\varepsilon\in(0,1/2)$ there is a ReLU network $f_{\mathsf{NN}}$ such that
The number of hidden layers $L$ of $f_{\mathsf{NN}}$ is bounded by $c\cdot\log\left(\varepsilon^{-n/(n-s)}\right)$ and the number of nonzero weights $M$ is bounded by $c\cdot\varepsilon^{-d/(n-s)}\cdot\log^{2}\left(\varepsilon^{-n/(n-s)}\right)$, where $c$ is a constant independent from $n$. Note that Gühring et al. [2020] refer to ReLU networks as defined in (16) as „standard neural networks“, since they also consider networks with skip connections.
Looking at the proof of Lemma 4.2 and using standard analysis, we can bound the second partial derivatives of $v^{n}_{t}$ by $\frac{C}{\sigma_{t}^{5}}$ for some $C>0$. Now we are ready to proof Theorem 3.6.
Proof.
From Lemma 4.1, we know that
$\mathsf{W}_{1}(\mathbb{P}^{\psi_{1}^{n}(Z)},\mathbb{P}^{\hat{\psi}_{1}(Z)})\leq\sqrt{2e}e^{\int_{0}^{1}\tilde{L}_{t}\;\mathrm{d}t}\inf_{\tilde{v}\in\mathcal{M}}\Big{(}\int_{0}^{1}\int_{(-a,a)^{d}}|v^{n}_{t}(\psi^{n}_{t}(x))-\tilde{v}_{t}(\psi^{n}_{t}(x))|^{2}p_{0}(x)\;\mathrm{d}x\;\mathrm{d}t+c\frac{a^{2d+2}e^{-da}}{\sigma_{\min}^{2}}\Big{)}^{1/2},$
where we used the choice of (10) before dividing the area in (18) in the proof of Lemma 4.1 and $\tilde{L}$ is the Lipschitz constant of $\tilde{v}$ at time $t$. We neglect an arbitrary small additional error on the right hand side to account for the use of the infimum, as the error bound corresponding to the kernel density estimator in (15) is nonzero anyway. Note that due to the restriction of the network class,
Integration yields for $n$ big enough
To bound the tail, we need to assure $a$ is chosen such that
As in the setting $d\geq 1,$ we can use that for $a\geq 15$
$a^{d+1}e^{-da/2}\leq e^{-\frac{da}{8}}.$
Now we want to bound the approximation term. We can bound
Note that $v_{t}^{n}\in C^{\infty}$ in case of the Gaussian kernel. Hence we can apply [Gühring et al., 2020, Corollary 4.2]. The bound of the second partial derivatives as well as the supremum norm bound of $v^{n}$ influence the coefficients in [Gühring et al., 2020, Corollary 4.2], therefore we construct the network such that $n^{-\frac{5}{2\alpha+d}}v^{n}$ is approximated and choose the last weight matrix such that it scales the result up. This addition of one layer does not influence the order of the layers and non zero weights of the network. However, the approximation error gets scaled up too, hence we need to ensure a smaller error. We need to choose $\varepsilon$ in (22) such that
The addition of one cutting layer to ensure the assumptions of Lemma 4.1 does not influence the order of the layers and non zero weights of the network either. Hence choosing
$L$ $\gtrsim\Big{(}\max\Big{(}\log(n)+n^{\frac{2}{2\alpha+d}},15\Big{)}\Big{)}^{d}\cdot\big{(}\log^{2}(n)+n^{\frac{4}{2\alpha+d}}\big{)},$
$M$ $\gtrsim\Big{(}\max\Big{(}\log(n)+n^{\frac{2}{2\alpha+d}},15\Big{)}\Big{)}^{d}\cdot\big{(}n^{\frac{12+2\alpha}{2\alpha+d}}e^{(4d+2)n^{\frac{2}{2\alpha+d}}}\big{)}^{d}\cdot\big{(}\log^{2}(n)+n^{\frac{4}{2\alpha+d}}\big{)},$
yields the desired rate. ∎
Proof of Theorem 3.8
Proof.
Recall the definitions $\mu_{n}\coloneqq\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$ and $K_{\sigma_{\min}}\coloneqq\frac{1}{\sigma_{\min}^{d}}K(\frac{\cdot}{\sigma_{\min}})$ from Theorem 3.3. We start again by decomposing
$\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ $\leq\sup_{f\in\mathrm{Lip}(1)}\int f(z)(p^{*}(z)-(K_{\sigma_{\min}}\ast p^{*})(z))\;\mathrm{d}z$
For the first term, we know from the proof of Theorem 3.3, using the Lebesgue measure restricted to $V$, that there is a $C_{1}$ such that
For the second term we use the orthogonality of $A$ to obtain
$=\sup_{f\in\mathrm{Lip}(1)}\int\frac{f(z)}{(2\pi\sigma_{\min}^{2})^{d/2}}\Big{(}\int\exp\Big{(}-\frac{\|z-x\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}(x)\;\mathrm{d}x-\frac{1}{n}\sum_{i=1}^{n}\exp\Big{(}-\frac{\|z-X_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int\frac{f(z)}{(2\pi\sigma_{\min}^{2})^{d/2}}\Big{(}\int\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|AA^{\top}z-x\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}(x)\;\mathrm{d}x$ (23)
$\qquad\qquad\qquad-\frac{1}{n}\sum_{i=1}^{n}\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|AA^{\top}z-X_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}z.$ (24)
By assumption, there is for every $x\in V$ a $y\in\mathbb{R}^{d^{\prime}}$ such that $x=Ay$. As the support of of $\mathbb{P}^{*}$ is subset of $V$, for every $X_{i}^{*}$ there is a $Y_{i}^{*}$ mapping into $\mathbb{R}^{d^{\prime}}$ such that $X_{i}^{*}=AY_{i}^{*}$ and $Y_{i}^{*}\sim p^{*}_{d^{\prime}}$. Hence we get
$\int\exp\Big{(}$ $-\frac{\|z-AA^{\top}z\|^{2}+\|AA^{\top}z-x\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}(x)\;\mathrm{d}x$
$=\int\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|AA^{\top}z-Ay\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y$
$=\int\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|A^{\top}z-y\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y,$
where the last equality holds due to the orthonormal columns of $A$. For (24) we proceed analogously. Now we can decompose every $z\in\mathbb{R}^{d}$ uniquely into
$z=Au+Bv,\quad\text{ for some }\quad u\in\mathbb{R}^{d^{\prime}},v\in\mathbb{R}^{d-d^{\prime}},$
where $B\in\mathbb{R}^{d\times(d-d^{\prime})}$ is a matrix that maps $\mathbb{R}^{d-d^{\prime}}$ into $V^{\perp}.$ As $A^{\top}B=0$ and therefore
$A^{\top}z=A^{\top}(Au+Bv)=A^{\top}Au=u\quad\text{ and similar }\quad(I_{d}-AA^{\top})z=v,$
we get for (23) and (24)
$\sup_{f\in\mathrm{Lip}(1)}\int\frac{f(z)}{(2\pi\sigma_{\min}^{2})^{d/2}}\Big{(}$ $\int\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|A^{\top}z-y\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y$
$-\frac{1}{n}\sum_{i=1}^{n}\exp\Big{(}-\frac{\|z-AA^{\top}z\|^{2}+\|A^{\top}z-Y_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}z$
$=\sup_{f\in\mathrm{Lip}(1)}\int\int\frac{f(Au+Bv)}{(2\pi\sigma_{\min}^{2})^{(d-d^{\prime})/2}}$ $\exp\Big{(}-\frac{\|v\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\;\mathrm{d}v\Big{(}\int\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-y\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y$
$-\frac{1}{n}\sum_{i=1}^{n}\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-Y_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}u.$
For fixed $v\in\mathbb{R}^{d-d^{\prime}}$ and $u,w\in\mathbb{R}^{(d-d^{\prime})}$, we have that
[…]
$\leq\int\frac{|f(Au+Bv)-f(Aw+Bv)|}{(2\pi\sigma_{\min}^{2})^{(d-d^{\prime})/2}}\exp\Big{(}-\frac{\|v\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\;\mathrm{d}v$
$\leq\int\frac{|A(u-w)|}{(2\pi\sigma_{\min}^{2})^{(d-d^{\prime})/2}}\exp\Big{(}-\frac{\|v\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\;\mathrm{d}v$
$=|u-w|,$
where we used the Lipschitz continuity of $f$ in the second inequality, the orthogonality of the columns of $A$ and the density of $\mathcal{N}(0,I_{d-d^{\prime}})$ in the last equality. We conclude that the integral is Lipschitz in $u$ with Lipschitz constant $1$. Hence
$\sup_{f\in\mathrm{Lip}(1)}\int\int\frac{f(Au+Bv)}{(2\pi\sigma_{\min}^{2})^{(d-d^{\prime})/2}}\exp\Big{(}-\frac{\|v\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\;\mathrm{d}v\Big{(}$ $\int\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-y\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y$
$-\frac{1}{n}\sum_{i=1}^{n}\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-Y_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}u$
$\leq\sup_{f\in\mathrm{Lip}(1)}\int f(u)\;\mathrm{d}v\Big{(}$ $\int\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-y\|^{2}}{2\sigma_{\min}^{2}}\Big{)}p^{*}_{d^{\prime}}(y)\;\mathrm{d}y$
$-\frac{1}{n}\sum_{i=1}^{n}\frac{1}{(2\pi\sigma_{\min}^{2})^{d^{\prime}/2}}\exp\Big{(}-\frac{\|u-Y_{i}^{*}\|^{2}}{2\sigma_{\min}^{2}}\Big{)}\Big{)}\;\mathrm{d}u.$
Now we can proceed along the proof of the second term of Theorem 3.3 to obtain the desired bound. ∎
Proof of Theorem 3.9
Proof.
We proceed completly analogeous to Theorem 3.6 replacing $d$ with $d^{\prime}$. However, since the vectorfield we want to approximate is still $d$ dimensional, we still need $\lceil 2a\rceil^{d}$ little cubes. Hence we choose
$L$ $\gtrsim\Big{(}\max\Big{(}\log(n)+n^{\frac{2}{2\alpha+d^{\prime}}},15\Big{)}\Big{)}^{d}\cdot\big{(}\log^{2}(n)+n^{\frac{4}{2\alpha+d^{\prime}}}\big{)},$
$M$ $\gtrsim\Big{(}\max\Big{(}\log(n)+n^{\frac{2}{2\alpha+d^{\prime}}},15\Big{)}\Big{)}^{d}\cdot\big{(}n^{\frac{12+2\alpha}{2\alpha+d^{\prime}}}e^{(4d^{\prime}+2)n^{\frac{2}{2\alpha+d^{\prime}}}}\big{)}^{d}\cdot\big{(}\log^{2}(n)+n^{\frac{4}{2\alpha+d^{\prime}}}\big{)},$
to obtain the desired rate. ∎

### v2

@@ 5.3 Auxiliary results
By assumption $\varphi$ decays faster than $e^{-x}.$ Using the upper incomplete $\Gamma$-function as defined in Lemma˜A.2, we obtain for $a>2$
@@
Proof of Lemma˜5.2.
@@
$=\frac{t}{\sigma_{t}^{2}}\Bigg(\frac{\sum_{i=1}^{n}X_{ik}X_{i\ell}p_{t}(x|X_{i})}{\sum_{j=1}^{n}p_{t}(x|X_{i})}-\frac{\big(\sum_{i=1}^{n}X_{ik}p_{t}(x|X_{i})\big)\big(\sum_{i=1}^{n}X_{i\ell}p_{t}(x|X_{i})\big)}{\big(\sum_{j=1}^{n}p_{t}(x|X_{i})\big)^{2}}\Bigg).$
@@
Proof of Lemma˜5.3.
Due to the triangle inequality, the largest increase in Euclidean distance is achieved when two points move in exact opposite direction. Since $\mathcal{M}$ has reach $\tau$ and $r<\tau$, the extreme case of this happens when $x$ and $y$ satisfy $|x-y|=2(\tau-r)$ and $|\pi(x)-\pi(y)|=2\tau$. This situation is sketched in Figure˜4. Note that the manifold does not need to "make a turn", the same reasoning applies to the bottleneck situation, for which we refer to Aamari et al. (2019, Figure 3).
Figure 4: Largest proportional displacement, black line is part of manifold $\mathcal{M}$, gray solid line is boundary of the reach environment, gray dotted lines are the boundaries of the $r$-tube around $\mathcal{M}$. Black dots are $x$ and $y$ respectively, the arrow symbolizes the maximal displacement.
Thus for all $x,y\in U_{r}$
$|\pi(x)-\pi(y)|\leq\frac{1}{1-\frac{r}{\tau}}|x-y|.\qed$
Proof of Lemma˜5.4.
  1. (I)
    Differentiating $g^{-1}(g(y))=y$ leads to $Dg^{-1}(g(y))Dg(y)=I_{d^{\prime}}$, which implies that $Dg^{-1}(g(y))$ is the pseudo-leftinverse of $Dg(y)$. Using the construction of the pseudoinverse via the singular value decomposition, we conclude that
    $\|Dg^{-1}(g(y))\|=\frac{1}{\lambda_{\min}(Dg(y))}\leq\frac{1}{m}<1.$
    Hence
    $\|A(y)\|\leq\|Dg^{-1}(g(y))\|\|P_{T_{g(y)}\mathcal{M}}\|\leq\frac{1}{m}.$
  2. (II)
    For any $h\in\mathbb{R}^{d^{\prime}}$, differentiating $Dg^{-1}(g(y))Dg(y)=I_{d^{\prime}}$ in direction $h$ yields
    $D(Dg^{-1}(g(y)))[h]Dg(y)+Dg^{-1}(g(y))D(Dg(y))[h]=0,$ (39)
    where $D(Dg(y))[h]=\frac{d}{dt}Dg(y+th)\big|_{t=0}$ and $D(Dg^{-1}(g(y)))[h]=\frac{d}{dt}Dg^{-1}(g(y+th))\big|_{t=0}$.
    Fix $y\in\mathbb{R}^{d^{\prime}}$ and set $x=g(y)$. Differentiating $g(g^{-1}(x))=x$ for yields
    $Dg(g^{-1}(x))Dg^{-1}(x)=I_{T_{x}\mathcal{M}}\quad\Longleftrightarrow\quad Dg(y)Dg^{-1}(g(y))=I_{T_{g(y)}\mathcal{M}},$
    where $I_{T_{x}\mathcal{M}}$ is the identity map on the tangent space. Using this right inverse of $Dg(y)$ and recalling that $P_{T_{g(y)}\mathcal{M}}$ is the orthogonal projection onto the tangent space, we obtain for (39),
    $D(Dg^{-1}(g(y)))[h]\Big|_{T_{g(y)}\mathcal{M}}=-Dg^{-1}(g(y))D(Dg(y))[h]Dg^{-1}(g(y)).$
    Taking the supremum over $h$ such that $|h|=1$ leads to
    $\|D(Dg^{-1}(g(y)))\|\leq\|Dg^{-1}(g(y))\|^{2}\|D(Dg(y))\|\leq\frac{M_{2}}{m^{2}}$
    Combining Stewart and Sun (1990, Chapter II, Theorem 4.5, Corollary 4.6) with Boissonnat et al. (2019, Corollary 3) and the fact that for $x\in[0,\frac{\pi}{2}]$ $\sin(x)\leq 3\sin\big(\frac{x}{2}\big)$ yields
    $\|P_{T_{g(y+th)}\mathcal{M}}-P_{T_{g(y)}\mathcal{M}}\|\leq\frac{3}{2\tau}|g(y+th)-g(y)|.$
    Furthermore,
    $\|D(P_{T_{g(y)}\mathcal{M}})[h]\|=\lim_{t\rightarrow 0}\frac{\|P_{T_{g(y+th)}\mathcal{M}}-P_{T_{g(y)}\mathcal{M}}\|}{|t|}\leq\lim_{t\rightarrow 0}\frac{3}{2\tau}\frac{|g(y+th)-g(y)|}{|t|}=\frac{3}{2\tau}|Dg(y)[h]|$
    Next, we differentiate for $h\in\mathbb{R}^{d^{\prime}}$
    $DA(y)[h]=D(Dg^{-1}(g(y)))[h]P_{T_{g(y)}\mathcal{M}}+Dg^{-1}(g(y))D(P_{T_{g(y)}\mathcal{M}})[h].$
    Recall that the operator norm is defined as $\sup_{|h|=1}\|DA(y)[h]\|=\sup_{|h|=1}\sup_{|v|=1}|DA(y)[h]v|$. Then, inserting all constants from ˜3.9
    $\|DA(y)\|\leq\|D(Dg^{-1}(g(y)))\|\|P_{T_{g(y)}\mathcal{M}}\|+\|Dg^{-1}(g(y))\|\|D(P_{T_{g(y)}\mathcal{M}})\|\leq\frac{M_{2}}{m^{2}}+\frac{3M_{1}}{2m\tau}=c_{0}<\frac{1}{r}.$
  3. (III)
    Assume that $F_{u}(y_{1})=F_{u}(y_{2})$. Then
    $0=F_{u}(y_{1})-F_{u}(y_{2})=y_{1}-y_{2}+(A(y_{1})-A(y_{2}))u.$
    If $y_{1}\neq y_{2}$
    $|y_{1}-y_{2}|=|(A(y_{1})-A(y_{2}))u|\leq|A(y_{1})-A(y_{2})||u|\leq c_{0}r|y_{1}-y_{2}|<|y_{1}-y_{2}|.$
    The above inequality is false, which implies $y_{1}=y_{2}$.
  4. (IV)
    Rearranging yields
    $F_{u}(S_{u}(z))=z\quad\Longleftrightarrow\quad S_{u}(z)+A(S_{u}(z))u=z\quad\Longleftrightarrow\quad S_{u}(z)=z-A(S_{u}(z)).$
  5. (V)
    In order to use the inverse function theorem, we first need to ensure that $DF_{u}$ has a nonzero determinant. For $h\in\mathbb{R}^{d^{\prime}}$, define
    $DF_{u}(y)[h]=h+(DA(y)[h])u.$
    Set the linear operator $G(y,u)\colon\mathbb{R}^{d^{\prime}}\rightarrow\mathbb{R}^{d^{\prime}}$ with $G(y,u)[h]=(DA(y)[h])u$, then
    $DF_{u}(y)=I_{d^{\prime}}+G(y,u),\quad\text{and}\quad\|G(y,u)\|\leq\|DA(y)\||u|\leq c_{0}r<1.$
    For the $i$-th largest singular value $s_{i}$, using Weyl’s singular value perturbation bound, for an English reference see e.g. Stewart (1990),
    $|s_{i}(DF_{u}(y))-s_{i}(I_{d^{\prime}})|\leq\|G(y,u)\|\leq\|DA(y)\||u|\leq c_{0}r<1.$
    Since the absolute value of the determinant is the product of the singular values, we obtain
    $(1-c_{0}r)^{d^{\prime}}\leq|\operatorname{det}(DF_{u}(y))|\leq(1+c_{0}r)^{d^{\prime}}.$
    Moreover, $t\mapsto\operatorname{det}\left(I_{d^{\prime}}+tG(y,u)\right)$ is continuous, unequal to 0 and equals 1 at $t=0$. Hence, $\operatorname{det}\left(DF_{u}(y)\right)>0$. Since $\|DA(y)\||u|\leq c_{0}r<1$, a Neumann series yields
    $\|(DF_{u}(y))^{-1}\|=\|(I_{d^{\prime}}-(-G(y,u)))^{-1}\|\leq(1-\|G(y,u)\|)^{-1}\leq\frac{1}{1-c_{0}r}$
    Since $F_{u}$ is injective, we have by the inverse function theorem that
    $DS_{u}(y)=(DF_{u}(S_{u}(y)))^{-1},\quad\text{ thus}\quad\|DS_{u}(y)\|\leq\frac{1}{1-c_{0}r}.$
  6. (VI)
    First, we differentiate $S_{u}(z)=z-A(S_{u}(z))u$ to obtain for $h\in\mathbb{R}^{d^{\prime}}$
    $DS_{u}(z)[h]=h-DA(S_{u}(z))[DS_{u}(z)[h]]u.$ (40)
    Defining the operator $B(z,u)\coloneqq(DA(S_{u}(z)))[\cdot]u$ for $v\in\mathbb{R}^{d^{\prime}}$ and rearranging yields on a matrix level
    $(I_{d^{\prime}}+B(z,u))DS_{u}(z)=I_{d^{\prime}}\quad\Longleftrightarrow\quad DS_{u}(z)=(I_{d^{\prime}}+B(z,u))^{-1},$
    where the invertibility follows from
    $\|B(z,u)\|=\sup_{|v|=1}\|(DA(S_{u}(z))[v])u\|\leq c_{0}|u|<1.$
    The inequality follows from property (II). Thus
    $\operatorname{det}(DS_{u}(z))=(\operatorname{det}(I+B(z,u)))^{-1}.$
    By the same singular value perturbation bound that was used in (V), we obtain
    $|s_{i}(I+B(z,u))-s_{i}(I)|\leq\|B(z,u)\|\leq c_{0}|u|.$
    This implies
    $(1-c_{0}|u|)^{d^{\prime}}\leq|\operatorname{det}(I+B(z,u))|\leq(1+c_{0}|u|)^{d^{\prime}},$
    which yields
    $(1+c_{0}|u|)^{-d^{\prime}}-1\leq|\operatorname{det}(DS_{u}(z))|-1\leq(1-c_{0}|u|)^{-d^{\prime}}-1.$
    We obtain
    $||\operatorname{det}(S_{u}(z))|-1|\leq\max\big(1-(1-c_{0}|u|)^{-d^{\prime}},(1+c_{0}|u|)^{-d^{\prime}}-1\big).$
    Two Taylor expansions yield for $\xi\in[0,c_{0}|u|]\subset[0,c_{0}r]$
    $|(1+\xi)^{-d^{\prime}}-1|$ $\leq\sup_{\zeta\in[0,c_{0}r]}d^{\prime}(1+\zeta)^{-d^{\prime}-1}\xi\leq d^{\prime}\xi,$
    $|(1-\xi)^{-d^{\prime}}-1|$ $\leq\sup_{\zeta\in[0,c_{0}r]}d^{\prime}(1-\zeta)^{-d^{\prime}-1}\xi\leq d^{\prime}(1-c_{0}r)^{d^{\prime}-1}\xi,$
    which implies
    $||\operatorname{det}(S_{u}(z))|-1|\leq d^{\prime}c_{0}(1-c_{0}r)^{d^{\prime}-1}|u|.$
Since $\sup_{y}\|DA(y)\||u|<1$, $F_{u}$ is surjective. Additionally, we showed that $S_{u}$ exists and from (40) and the properties of $g$ we can conclude that the derivative of $S_{u}$ is continuous, we obtain that $F_{u}$ is a diffeomorphism. ∎